# When Truth Is Overridden: Uncovering the Internal Origins of Sycophancy in Large Language Models

### v2

@@ Conclusion
This study provides a mechanistic explanation for sycophancy in LLMs, showing it is opinion-driven, not authority-driven. Models consistently agree with incorrect user opinions regardless of their claimed expertise, as they fail to represent authority internally. In contrast, user opinions actively suppress the model’s learned knowledge in later layers, validated by causal activation patching. The study also identifies a powerful perspective-driven effect: first-person prompts (“I believe…”) induce significantly more sycophancy than third-person ones (“They believe…”) by creating a stronger override of the model’s internal knowledge.
@@ Appendix A Appendix
The logit-lens, as described by nostalgebraist (2020) , provides a way to “decode” the model’s intermediate representations by projecting them through the final language modeling head at each layer. This allows us to inspect what each layer “believes” the next token should be, even before the final output. Notably, earlier layers often produce distributions that are less peaked and more generic, while later layers refine these predictions, sometimes correcting earlier mistakes or amplifying certain hypotheses. This technique can reveal how information and uncertainty propagate through the network, and can highlight when and where the model commits to specific predictions. Specifically, for a model with $L$ layers, the probability distribution over the next token at layer $l$ is given by:
@@
We used GPT-4o (Hurst et al. 2024) to obtain the prefix library in three steps.
@@
Additional Experimentalal Results

### v3

@@ Conclusion
This study offers a mechanistic explanation for sycophancy in LLMs, showing it is opinion- rather than authority-driven, as they fail to represent authority internally. User opinions suppress learned knowledge in later layers, validated by causal activation patching. We also identify a strong perspective-driven effect: first-person prompts elicit more sycophancy than third-person ones by creating a stronger override of the model’s internal knowledge.
Acknowledgements
Di Wang and Shu Yang are supported in part by the funding BAS/1/1689-01-01 and funding from KAUST - Center of Excellence for Generative AI, under award number 5940 and a gift from Google.
@@ Appendix A Appendix
The logit-lens, as described by nostalgebraist2020interpreting , provides a way to “decode” the model’s intermediate representations by projecting them through the final language modeling head at each layer. This allows us to inspect what each layer “believes” the next token should be, even before the final output. Notably, earlier layers often produce distributions that are less peaked and more generic, while later layers refine these predictions, sometimes correcting earlier mistakes or amplifying certain hypotheses. This technique can reveal how information and uncertainty propagate through the network, and can highlight when and where the model commits to specific predictions. Specifically, for a model with $L$ layers, the probability distribution over the next token at layer $l$ is given by:
@@
We used GPT-4o (hurst2024gpt) to obtain the prefix library in three steps.
@@
Additional Experimental Results